OpenStudy (anonymous):

At the motocross races, three hamburgers and six soft drinks cost $24, and three hamburgers and three soft drinks cost $18. Which system of linear equations below can be used to determine the price of a hamburger and the price of a soft drink? Let x represent the cost of a hamburger, and let y represent the cost of a soft drink. 3x + 6y = 24 - a 3x + 3y = 18 3x + 3y = 24 - b 6x + 3y = 18 6x + 9y = 42 - c 3x + 3x = 24 - d 6y + 3y = 18

OpenStudy (anonymous):

Need Help @ARMYRANGERS

OpenStudy (nurali):

let x = the price of one hamburger y = the price of one soft drink three hamburgers and six soft drinks cost $24 3x + 6y = 24 three hamburgers and three soft drinks cost $18 3x + 3y = 18 the system of linear equations used to determine the price is 3x + 6y = 24 3x + 3y = 18

OpenStudy (anonymous):

Thanks

OpenStudy (nurali):

My Pleasure.

OpenStudy (anonymous):

Do you work at truetts or chickfila